1) The sum of two numbers is 89. Their difference is 13. What are the two numbers?
Pls indicate the representation,the equation,solution,checking and the cheking.
Thank you!

Found 2 solutions by mananth, rfadrogane:
Answer by mananth(16949)

let the numbers be x & y
..
x+y =89
x-y =13
...
add the equations
2x=89+13
2x=102
/2
x=51
..
x+y = 89
x=51
y=89-51
y=38
...
51 & 38 are the numbers
..
m.ananth@hotmail.ca

Answer by rfadrogane(214)

The sum of two numbers is 89. Their difference is 13. What are the two numbers?
Sol'n:
let x & y - the numbers
x+y = 89 ---(1)
x-y = 13 ----(2)
Add, the 2 equations
2x = 102, divide by 2 on both side
x = 51
for y:
y = 89-x
= 89-51
y = 38
thus:
x = 51 & y = 38 ----answer
rfadrogane,ree-rme